OpenStudy (anonymous):

PLEASE HELP write two limit statements that describe the end behavior of y=(x^4-2x^3+x^2)/(3x^3-9x^2+6x)

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

i think there are four limit statements since there are two graphs

OpenStudy (anonymous):

y=(x^4-2x^3+x^2)/(3x^3-9x^2+6x)

OpenStudy (anonymous):

as \(x\to \infty, f(x)\to \infty\)

OpenStudy (anonymous):

y=(x^4-2x^3+x^2) / (3x^3-9x^2+6x) y=(x^3-2x^2+x) / (3x^2-9x+6) y=(x^3-2x^2+x) / (3x^2-9x+6) y=x(x^2-2x+1) / 3(x^2-3x+2) y=x(x-1)^2 / 3(x-1)(x-2) y=x(x-1) / 3(x-1) Perhaps that simplifies that a little.

OpenStudy (anonymous):

I mean y=x(x-1) / 3(x-2)

OpenStudy (anonymous):

yea i got that part

OpenStudy (anonymous):

if you divide you get a polynomial of degree 1 (a line) with positive slope and that tells you that as \(x\to \infty,\) then \(y\to \infty\) and if \(x\to -\infty\) then \(y\to -\infty\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay so i have a question. we simplified that to f(x) = x(x-1)/3(x-2) so what are the domain restrictions

OpenStudy (anonymous):

well 3(x-2) can't be zero, so x can't equal what ?

OpenStudy (anonymous):

the domain restrictions are computed BEFORE you factor and cancel

OpenStudy (anonymous):

the domain restrictions for the entire graph is all real numbers except 0, 2, and 1

OpenStudy (anonymous):

ma bad

OpenStudy (anonymous):

but the worksheet says to find the domain restrictions based on the simplified equation

OpenStudy (anonymous):

then it was written by a moron

OpenStudy (anonymous):

haha okay

OpenStudy (anonymous):

you cannot cancel zeros

OpenStudy (anonymous):

so the vertical asymptote is at x = 2

OpenStudy (anonymous):

so the limits are as follows right?

OpenStudy (anonymous):

vertical asymptote x = 2 right

OpenStudy (anonymous):